# SPRMamba: Surgical Phase Recognition for Endoscopic Submucosal Dissection with Mamba

### v1

@@ 4 Experiments
4.1 Dataset
| ESD Surgical phases | Preparation | Estimation | Marking | Injection | Incision | ESD | Vessel-treatment | Clips | Total |
| Train | 91739 | 43198 | 15585 | 44464 | 59695 | 174201 | 44317 | 10817 | 484016 |
| val | 14193 | 5822 | 2820 | 9886 | 11917 | 41807 | 4865 | 3985 | 95295 |
| test | 14323 | 16860 | 6583 | 15831 | 22181 | 75968 | 14707 | 4748 | 171201 |
4.1.1 ESD385
@@
There is no publicly available dataset on ESD surgery. To validate the effectiveness of the proposed algorithm in this study, we retrospectively selected patients who underwent ESD from August 16, 2023, to January 8, 2024, in the endoscopy unit of the Department of Gastroenterology, Renji Hospital, Shanghai Jiao Tong University. A total of 385 videos of ESD procedures were collected. All procedures were performed by experienced endoscopists. The instruments used for ESD procedures included: gastroscope GIF-Q260, GIF-H260, GIF-HQ290, mucosal incision knife KD-612L/U, KD-655L/U, KD-620UR, endoscopic water pump OFP-2, mucosal injection needle NM-400L-0423, NM-400U-0423, thermal biopsy forceps HDBF-2.4-230-S, hemostatic forceps FD-410LR, FD-412LR, soft tissue clips ROCC-D-26-195-C, ROCC-D-26-235-C, hemostatic clips M00521242, methylene blue injection, indigo rouge solution. All endoscopic surgical videos were recorded using Image Management Hub, IMH-200, and Olympus and stored in MP4 format. The video resolution was 1920 × 1080 with a frame rate of 50 fps. After collection, all videos were anonymized to remove identifiable patient information, procedure timestamps, and other identifying details. This work was approved by the Renji Hospital affiliated with the Shanghai Jiao Tong University School of Medicine ethics committee with number RA-2024-457(2024.3.20).
@@
The ESD video is divided into eight surgical phases, including (1) Estimation; (2) Marking; (3) Injection; (4) Incision; (5) ESD; (6) Vessel-treatment; (7) Clips(portion); and (8) Preparation. Phase (6) Preparation included elements unrelated to the ESD procedure, such as gastric insufflation and device replacement. Four endoscopists independently annotated the video, labeling each frame in the video. After the initial annotation, quality control was performed by two additional experienced endoscopists. Any uncertainties that arose during the quality control process were resolved through collaborative discussions among the six experts. The number of annotated frames in the dataset varied at each phase, with the ESD phase occupying most of the procedure time, which is the most important and skill-demanding phase of ESD. Detailed statistics for each phase are shown in Table 7. Overall, a total of 484016 frames, 95,295 and 171,201 frames were annotated for training, validation, and verification, respectively.
4.1.2 Cholec80
@@
Furthermore, to verify the robustness of our proposed algorithm, we validated it on the Cholec80 dataset[3] for cholecystectomy procedures. The Cholec80 dataset is a large-scale surgical benchmark dataset containing 80 cholecystectomy videos performed by 13 surgeons. These videos are recorded at 25 frames per second, with each frame having a resolution of either 1920×1080 or 854×480 pixels. Each frame in the videos is annotated with one of seven surgical phases: Preparation (P1), Calot triangle dissection (P2), Clipping and cutting (P3), Gallbladder dissection (P4), Gallbladder packaging (P5), Cleaning and coagulation (P6), and Gallbladder retraction (P7). Additionally, each frame includes annotations for seven types of surgical tools, including Grasper, Bipolar, Hook, Clipper, Scissors, Irrigator, and Specimen bag. We strictly followed the experimental protocol used in previous studies[3, 26, 6], dividing the dataset into two equal subsets: the first 40 videos for training and the remaining 40 videos for testing, with 8 videos from the training set selected for validation.
4.2 Evaluation Metrics
@@
For surgical phase recognition, we employed four commonly used evaluation metrics to quantitatively assess model performance, which have been used in previous phase recognition work as well[3, 26, 45, 6]. These metrics include Precision, Recall, Jaccard Index, and Accuracy. Accuracy is defined as the percentage of frames across the entire video correctly predicted to be in their ground truth phase. Given the imbalanced phase presented in the video, Precision, Recall, and Jaccard Index refer to phase-level evaluations, calculated within each phase and then averaged across all phases.
4.3 Implementation details
@@
Our approach is implemented in Python using Pytorch framework and training is conducted on a workstation equipped with an NVIDIA RTX 4090. In the first stage, we use a ResNet-50 model pre-trained on ImageNet-22K[44]. We then fine-tune the model on our data. To ensure a fair comparison with SOTA methods, we downsample all videos to 1 fps, which is also the approach used in previous works[45, 6]. This operation has additional benefits, including enriching temporal information and saving memory. During training, image frames are further downsampled from the original resolutions of 1920 1080 and 854 480 to a resolution of 250 250 pixels to further reduce memory usage. Data augmentation is performed through 224 224 cropping, flipping, and random mirroring to expand the training dataset. The ResNet-50 model is fine-tuned with a batch size of 160 images. In the second stage, LSTContext is trained end-to-end from scratch. The W and G for each task in the LSTContext module are initially chosen based on estimates of task average durations on the training dataset and are then fine-tuned through ablation studies. For both stages, we adopted the same experimental setup as the literature[44], using the AdamW optimizer[46] with a weight decay of 1e-5 and approximately 200 iterations. A cosine learning rate scheduler[47] is used, with 40 epochs of linear warm-up and an initial learning rate of 5e-4.
@@ 4.4.1 Result on the ESD385 Dataset
| SPRMamba(Ours) | 87.649.83 | 86.7216.54 | 86.7615.66 | 77.5117.83 | 7.5G | 25.42M |
On the ESD385 dataset, we compared our proposed method with the SOTA method. This includes the method proposed by Furube et al. [12], which fine-tunes ResNet-50 as a feature extractor and then uses MS-TCN for hierarchical prediction refinement for surgical phase recognition. The method proposed by Cao et al. [11]. An Intelligent Surgical Workflow Recognition Suite for ESD is based on Trans-SVNet [9] implementation. In addition, we compared our method with SV-RCNet [26] and SAHC [48], two SOTA methods for cholecystectomy surgical phase recognition. The comparative results are presented in Table 2. We found that applying existing surgical phase recognition methods directly to ESD leads to performance degradation because ESD surgery has more complex workflows compared to traditional surgeries. In the case of image classification using ResNet-50 alone, the average accuracy is 72.31%, the average precision is 78.75%, the average recall is 66.63%, and the average Jaccard is 55.39%. After including the SRTM module and Temporal Sample Strategy, the present method achieves an average accuracy of 87.64%, an average precision of 86.72%, an average recall of 86.76%, and an average Jaccard of 77.51%, increasing the average accuracy from 72.31% to 87.64%. These results demonstrate the effectiveness of the method for automatic phase recognition. Compared to the SOTA method, our method outperforms the second-best method in all four metrics by 1.00%, 0.37%, 3.01%, and 2.37%, respectively. Furthermore, to demonstrate the algorithm efficiency of the proposed method. We also compare the number of parameters and computational cost of our method with the SOTA method, which is reported in Table 2. "Params" denotes the number of parameters. Our proposed method outperforms SAHC in terms of fewer parameters. Compared to SAHC, our model will reduce 2.0G FLOPS.
In Fig. 3, we visualized the predictions of two ESD videos to qualitatively show the improvement in surgical phase recognition. The results highlight that SPRMamba obtains consistent and smooth predictions not only within one phase but also for the often ambiguous phase transitions. Compared with AI-Endo and Furube et at., SPRMamba can perform accurate phase recognition even for phases with short durations, such as phase Preparation and Clips. Finally, SPRMamba shows robustness because Video 2 lacks phase Clips. However, the performance of our model does not deteriorate.
@@ 4.4.2 Result on the Cholec80 Dataset
In the Cholec80 dataset, we conducted a performance comparison of our methods with state-of-the-art (SOTA) methods, including Endonet [3], MTRCNet-CL [45], LAST [49], SV-RCNet [26], TeCNO [7], TMRNet [6], Trans-SVNet [9], and SAHC [48]. Please note that we re-implemented TeCNO and Trans-SVNet using the model weights provided in the original manuscript. The results of the other state-of-the-art methods were extracted verbatim from their respective published works. Our comparison results are presented in Table 3. Our method outperforms the other methods in most of the evaluation metrics, except for the average precision, where it ranks third behind MTRCNet-CL and SAHC. Specifically, for average accuracy, SPRMamba matches the high accuracy of LAST (a multi-task learning method that requires additional information in the form of extra labels) but displays a lower standard deviation of approximately 0.13%. Additionally, for average recall and average Jaccard, SPRMamba achieves the best performance with 90.12% and 81.43%, respectively.
Furthermore, to illustrate the performance of our approach compared to state-of-the-art methods, in Fig. 4 we qualitatively compare two examples from the Cholec80 testing dataset. We compare the proposed method with two SOTA methods, TeCNO [7] and Trans-SVNet [9]. As shown in Fig. 4, our method predicts higher frame-level accuracy, especially on the boundary between two different phases. In addition, we can notice that Trans-SVNet and TeCNO make a lot of mistakes, i.e., they are unable to accurately classify ambiguous frames in different phases. For example, some frames within P3 are misclassified into P4 and some frames within P5 are misclassified into P4 in Video 1. On the contrary, our proposed method performs very accurately compared to ground truth values.
@@ 4.5.1 Scale Residual TranMamba
| Accuracy(%) | Precision(%) | Recall(%) | Jaccard(%) | FLOPs | Params | |
| Transformer | 86.0310.89 | 86.0217.40 | 84.8917.03 | 75.4418.41 | 8.5G | 26.15M |
| Mamba | 86.9112.06 | 86.5917.24 | 86.0516.74 | 77.1918.42 | 6.5G | 25.28M |
| SRTM | 87.649.83 | 86.7216.54 | 86.7615.66 | 77.5117.83 | 7.5G | 25.42M |
@@
4.5.2 Temporal Sample Strategy
@@
| Ours | 87.649.83 | 86.7216.54 | 86.7615.66 | 77.5117.83 |
To demonstrate the effectiveness of the Temporal Sample Strategy, we designed four ablation experiments. We also configured the original SRTM framework matching the SPRMamba structure as a benchmark for comparison. As shown in Table 5, baseline achieves an average accuracy of 86.58%, an average precision of 86.60%, an average recall of 84.16%, and an average Jaccard of 75.65% for surgical phase recognition. The experimental results outperform the baseline modules regardless of whether the key modules are added individually or in combination with each other. First, we show the impact of combining STContext SRTM and LTContext SRTM in the LSTContext block shown in Fig 2. We compare it with two variants where only STContext SRTM or LTContext SRTM is used. To keep the number of parameters constant, in these cases, we still use both SRTM blocks in LSTContext. The results show that combining the two SRTM blocks gives better results.
@@ 4.5.3 Effect of Different value of W and G
To further explore the effect of the Temporal Sample Strategy with different W and G on the model performance, we conducted the experiments shown in Table 6. The parameter W controls the size of the local window and the parameter G controls the range of the long-term temporal context modeling. The experimental results show that the performance of the models (Accuracy and Jaccard) increases as G increases until the optimal performance is reached for G=64. The performance fluctuates as the size of the local window W becomes smaller, but W=64 works best.

### v2

@@ 4 Experiments and Results Analysis
4.1 ESD385 Dataset
@@
ESD385 retrospectively selected patients who underwent ESD from August 16, 2023, to January 8, 2024, in the endoscopy unit of the Department of Gastroenterology, Renji Hospital, Shanghai Jiao Tong University. A total of 385 videos of ESD procedures were collected. All procedures were performed by experienced endoscopists. The instruments used for ESD procedures included: gastroscope GIF-Q260, GIF-H260, GIF-HQ290, mucosal incision knife KD-612L/U, KD-655L/U, KD-620UR, endoscopic water pump OFP-2, mucosal injection needle NM-400L-0423, NM-400U-0423, thermal biopsy forceps HDBF-2.4-230-S, hemostatic forceps FD-410LR, FD-412LR, soft tissue clips ROCC-D-26-195-C, ROCC-D-26-235-C, hemostatic clips M00521242, methylene blue injection, indigo rouge solution. All endoscopic surgical videos were recorded using Image Management Hub, IMH-200, and Olympus and stored in MP4 format. The video resolution was 1920 × 1080 with a frame rate of 50 fps. After collection, all videos were anonymized to remove identifiable patient information, procedure timestamps, and other identifying details. This work was approved by the Renji Hospital affiliated with the Shanghai Jiao Tong University School of Medicine ethics committee with number RA-2024-457(2024.3.20).
| Train | Val | Test | |
| Preparation | 91739 | 14193 | 14323 |
| Estimation | 43198 | 5822 | 16860 |
| Marking | 15585 | 2820 | 6583 |
| Injection | 44464 | 9886 | 15831 |
| Incision | 59695 | 11917 | 22181 |
| ESD | 174201 | 41807 | 75968 |
| Vessel-treatment | 44317 | 4865 | 14707 |
| Clips | 10817 | 3985 | 4748 |
| Total | 484016 | 95295 | 171201 |
@@
The ESD video is divided into eight surgical phases, including (1) Preparation; (2) Estimation; (3) Marking; (4) Injection; (5) Incision; (6) ESD; (7) Vessel-treatment; and (8) Clips(portion). Samples are shown in Fig.5. Phase (1) Preparation included elements unrelated to the ESD procedure, such as gastric insufflation and device replacement. Four endoscopists independently annotated the video, labeling each frame in the video. After the initial annotation, quality control was performed by two additional experienced endoscopists. Any uncertainties that arose during the quality control process were resolved through collaborative discussions among the six experts. The number of annotated frames in the dataset varied at each phase, with the ESD phase occupying most of the procedure time, which is the most important and skill-demanding phase of ESD. Detailed statistics for each phase are shown in Table 9. Overall, a total of 484016 frames, 95295, and 171201 frames were annotated for training, validation, and verification, respectively.
@@
Cholec80 is a large-scale surgical benchmark dataset containing 80 cholecystectomy videos performed by 13 surgeons. These videos are recorded at 25 frames per second, with each frame having a resolution of either 1920×1080 or 854×480 pixels. Each frame in the videos is annotated with one of seven surgical phases: Preparation (P1), Calot triangle dissection (P2), Clipping and cutting (P3), Gallbladder dissection (P4), Gallbladder packaging (P5), Cleaning and coagulation (P6), and Gallbladder retraction (P7). Additionally, each frame includes annotations for seven types of surgical tools, including Grasper, Bipolar, Hook, Clipper, Scissors, Irrigator, and Specimen bag. We strictly followed the experimental protocol used in previous studies [3, 26, 6], dividing the dataset into two equal subsets: the first 40 videos for training and the remaining 40 videos for testing, with 8 videos from the training set selected for validation.
4.2 Experiments Design
@@
Evalutation Metrics. For surgical phase recognition, we employed four commonly used evaluation metrics to quantitatively assess model performance, which have been used in previous phase recognition work as well [3, 26, 45, 6]. These metrics include Precision, Recall, Jaccard Index, and Accuracy. Accuracy is defined as the percentage of frames across the entire video correctly predicted to be in their ground truth phase. Given the imbalanced phase presented in the video, Precision, Recall, and Jaccard Index refer to phase-level evaluations, calculated within each phase and then averaged across all phases.
@@
Implementation details. Our approach is implemented in Python using the PyTorch framework, and training is conducted on a workstation equipped with an NVIDIA RTX 4090. In the first stage, we use a ResNet-50 model pre-trained on ImageNet-22K [44]. We then fine-tune the model on our data. To ensure a fair comparison with SOTA methods, we downsample all videos to 1 fps, which is also the approach used in previous works [45, 6]. This operation has additional benefits, including enriching temporal information and saving memory. During training, image frames are further downsampled from the original resolutions of 1920 1080 and 854 480 to a resolution of 250 250 pixels to further reduce memory usage. Data augmentation is performed through 224 224 cropping, flipping, and random mirroring to expand the training dataset. The ResNet-50 model is fine-tuned with a batch size of 160 images. In the second stage, LSTContext is trained end-to-end from scratch. The W and G for each task in the LSTContext module are initially chosen based on estimates of task average durations on the training dataset and are then fine-tuned through ablation studies. For both stages, we adopted the same experimental setup as the literature [44], using the AdamW optimizer [46] with a weight decay of 1e-5 and approximately 200 iterations. A cosine learning rate scheduler [47] is used, with 40 epochs of linear warm-up and an initial learning rate of 5e-4.
@@ 4.3.1 Result on the ESD385 Dataset
On the ESD385 dataset, we compared our proposed method with the SOTA method. This includes the method proposed by Furube et al. [12], which fine-tunes ResNet-50 as a feature extractor and then uses MS-TCN for hierarchical prediction refinement for surgical phase recognition. The method proposed by Cao et al. [11]. An Intelligent Surgical Workflow Recognition Suite for ESD is based on Trans-SVNet [9] implementation. In addition, we compared our method with SV-RCNet [26] and SAHC [48], two SOTA methods for cholecystectomy surgical phase recognition.
@@
| SPRMamba(Ours) | 87.7011.03 | 86.8416.07 | 86.8915.64 | 77.7417.60 | 7.5G | 25.42M |
The comparative results are presented in Table 2. We found that applying existing surgical phase recognition methods directly to ESD leads to performance degradation because ESD surgery has more complex workflows compared to traditional surgeries. In the case of image classification using ResNet-50 alone, the average accuracy is 72.31%, the average precision is 78.75%, the average recall is 66.63%, and the average Jaccard is 55.39%. After including the SRTM module and Hierarchical Sampling Strateg, the present method achieves an average accuracy of 87.64%, an average precision of 86.72%, an average recall of 86.76%, and an average Jaccard of 77.51%, increasing the average accuracy from 72.31% to 87.64%. These results demonstrate the effectiveness of the method for automatic phase recognition. Compared to the SOTA method, our method outperforms the second-best method in all four metrics by 1.00%, 0.37%, 3.01%, and 2.37%, respectively. Furthermore, to demonstrate the algorithm efficiency of the proposed method. We also compare the number of parameters and computational cost of our method with the SOTA method, which is reported in Table 2. "Params" denotes the number of parameters. Our proposed method outperforms SAHC in terms of fewer parameters. Compared to SAHC, our model will reduce 2.0G FLOPS.
In Fig.6, we visualized the predictions of two ESD videos to qualitatively show the improvement in surgical phase recognition. The results highlight that SPRMamba obtains consistent and smooth predictions not only within one phase but also for the often ambiguous phase transitions. Compared with AI-Endo and Furube et at., SPRMamba can perform accurate phase recognition even for phases with short durations, such as phase Preparation and Clips. Finally, SPRMamba shows robustness because case 2 lacks phase Clips. However, the performance of our model does not deteriorate.
@@ 4.3.2 Result on the Cholec80 Dataset
In the Cholec80 dataset, we conducted a performance comparison of our methods with state-of-the-art (SOTA) methods, including Endonet [3], MTRCNet-CL [45], LAST [49], SV-RCNet [26], TeCNO [7], TMRNet [6], Trans-SVNet [9], and SAHC [48]. Please note that we re-implemented TeCNO and Trans-SVNet using the model weights provided in the original manuscript. The results of the other state-of-the-art methods were extracted verbatim from their respective published works.
@@
Our comparison results are presented in Table 3. Our method outperforms the other methods in most of the evaluation metrics, except for the average precision, where it ranks third behind MTRCNet-CL and SAHC. Specifically, for average accuracy, SPRMamba matches the high accuracy of LAST (a multi-task learning method that requires additional information in the form of extra labels) but displays a lower standard deviation of approximately 0.13%. Additionally, for average recall and average Jaccard, SPRMamba achieves the best performance with 90.12% and 81.43%, respectively.
Furthermore, to further demonstrate the effectiveness of our approach, Fig.7 presents a qualitative comparison between SPRMamba and two state-of-the-art methods (TeCNO [7] and Trans-SVNet [9]) on representative samples from the Cholec80 test set. The results show that SPRMamba achieves more accurate phase recognition, particularly at transition boundaries where competing methods frequently misclassify frames, such as between phases P3/P4 and P5/P4. While TeCNO and Trans-SVNet exhibit noticeable instability with scattered misclassifications, our method maintains strong alignment with ground truth labels, reflecting its robustness in handling ambiguous frames and preserving temporal coherence. This improvement stems from SPRMamba’s ability to effectively model spatial-temporal dependencies, reducing both boundary errors and over-segmentation artifacts observed in existing approaches. The visual analysis corroborates our quantitative findings, further validating the advantages of the proposed architecture.
@@ 4.4.1 Scale Residual TranMamba
| Accuracy(%) | Precision(%) | Recall(%) | Jaccard(%) | |
| Transformer | 86.0310.89 | 86.0217.40 | 84.8917.03 | 75.4418.41 |
| Mamba | 86.9112.06 | 86.5917.24 | 86.0516.74 | 77.1918.42 |
| SRTM(Ours) | 87.7011.03 | 86.8416.07 | 86.8915.64 | 77.7417.60 |
@@
4.4.2 Hierarchical Sampling Strategy
@@
| SPRMamba(Ours) | 87.7011.03 | 86.8416.07 | 86.8915.64 | 77.7417.60 |
To demonstrate the effectiveness of the Hierarchical Sampling Strategy, we designed four ablation experiments. We also configured the original SRTM framework matching the SPRMamba structure as a benchmark for comparison. As shown in Table 5, baseline achieves an average accuracy of 86.58%, an average precision of 86.60%, an average recall of 84.16%, and an average Jaccard of 75.65% for surgical phase recognition. The experimental results outperform the baseline modules regardless of whether the key modules are added individually or in combination with each other. First, we show the impact of combining STContext SRTM and LTContext SRTM in the LSTContext block shown in Fig 2. We compare it with two variants where only STContext SRTM or LTContext SRTM is used. To keep the number of parameters constant, in these cases, we still use both SRTM blocks in LSTContext. The results show that combining the two SRTM blocks gives better results.
4.4.3 Effect of Different Number Of Layers
| Layers | Accuracy(%) | Precision(%) | Recall(%) | Jaccard(%) |
| 8 | 86.9111.17 | 86.8817.37 | 84.8717.28 | 76.3118.16 |
| 9 | 86.9410.78 | 87.3717.82 | 84.9616.93 | 76.7617.77 |
| 10 | 87.7011.03 | 86.8416.07 | 86.8915.64 | 77.7417.60 |
| 11 | 87.5010.75 | 87.4917.04 | 85.6817.37 | 77.4118.40 |
| 12 | 87.649.83 | 87.1717.49 | 85.4916.01 | 77.1117.77 |
The ablation study investigating the impact of varying the number of layers in the LSTContext blocks is presented in Table 6. The results demonstrate that the model’s performance is generally robust to changes in layer depth, with all configurations achieving competitive metrics. Notably, the 10-layer configuration achieved the highest accuracy, recall, and Jaccard, while the 11-layer variant yielded the best precision. The marginal differences between the 10-, 11-, and 12-layer models suggest that the architecture saturates in performance beyond 10 layers, with diminishing returns for additional depth. Based on these findings, we selected N=10 for the ESD385 datasets, respectively, balancing computational efficiency and task-specific requirements.
4.4.4 Effect of Different Number Of Stages
| Stages | Accuracy(%) | Precision(%) | Recall(%) | Jaccard(%) |
| 1 | 87.1611.07 | 86.8617.51 | 84.8518.18 | 76.3618.79 |
| 2 | 86.6811.22 | 86.7717.45 | 85.1617.12 | 76.4318.38 |
| 3 | 87.3110.40 | 86.8518.17 | 85.0117.03 | 76.5118.18 |
| 4 | 87.7011.03 | 86.8416.07 | 86.8915.64 | 77.7417.60 |
| 5 | 87.1410.26 | 87.1816.57 | 85.6616.01 | 77.2016.64 |
As shown in Fig.2, we use four stages of LSTContext blocks, each of them with 10 layers. To assess the influence of stage depth, we conducted an ablation study, with the results summarized in Table 7. The analysis reveals that employing multiple stages effectively mitigates over-segmentation errors while substantially enhancing both the Accuracy and Jaccard compared to using only single-stage. Performance metrics consistently improve as the number of stages increases up to four, achieving optimal accuracy, recall, and Jaccard. However, expanding to five stages leads to a decline in performance, indicating the onset of overfitting. These findings validate the selection of four stages as the optimal balance between model complexity and generalization capability.
@@ 4.4.5 Effect of Different Value of W and G
To further explore the effect of the Hierarchical Sampling Strategy with different W and G on the model performance, we conducted the experiments shown in Table 8. The parameter W controls the size of the local window and the parameter G controls the range of the long-term temporal context modeling. The experimental results show that the performance of the models (Accuracy and Jaccard) increases as G increases until the optimal performance is reached for G=64. The performance fluctuates as the size of the local window W becomes smaller, but W=64 works best.